# IndiSeek learns information-guided disentangled representations

### v3

@@ 1 Introduction
In contrast, identifying modality-specific features, also known as disentangled representation learning, has become an active research area (fischer2020conditional; liang2023factorized; liu2023focal; wang2024information; dufumier2024align; wang2024decoupling). While shared features should be minimally sufficient, modality-specific features must both remain independent of (i.e., disentangled from) shared features and capture all complementary information within each modality (liang2023factorized; wang2024information; wang2024decoupling). From an information-theoretic perspective, this requires (1) simultaneously minimizing the mutual information between shared and modality-specific features and (2) maximizing the mutual information between data and their union. Multiple state-of-the-art methods (liang2023factorized; wang2024information) tackle this problem from such a perspective, but differ in how they approximate the underlying mutual information. Since such quantities are difficult to estimate with finite samples, these methods often trade off independence against completeness. As we show in the motivating examples below, this tradeoff limits their ability to achieve both goals simultaneously.
@@ 1.1 Motivating examples and limitations of SOTA
We present the average coordinate importance metrics in learned modality-specific representation maps over $50$ simulation runs in Figures 1 and 2 for three different learning algorithms in the two settings, respectively. In both figures, the left panels report results of the SOTA approach FactorizedCL in liang2023factorized (without self-supervision), the middle panels report results of the SOTA approach InfoDisen in wang2024information, and the right panels report results of the new approach IndiSeek we propose in this paper. Tuning parameters [^2] See Appendix A.2 for an ablation study on performances with different tuning parameter values. are set at $0.01$ for InfoDisen and IndiSeek and at $1$ by default for FactorizedCL.
As we have reasoned when introducing the settings, the ideal modality-specific representations in Setting 1 should depend on and only on the last four coordinates, while in Setting 2, on and only the last two coordinates. The left and the middle panels of both figures suggest that neither FactorizedCL nor InfoDisen learn an ideal representation in either setting, as irrelevant coordinates play non-trivial roles in determining the learned modality-specific representations. Therefore, both SOTA approaches have their limitations in achieving the desired goal of disentangled representation learning. In contrast, the right panels of Figures 1 and 2 show that the new method IndiSeek we are to propose achieves it in both settings: the learned modality-specific features depend on and only on all coordinates that are independent of those involved in shared features.
@@ 2 Ideal decomposition of multi-modal data
    Independence: modality-specific and shared features should be disentangled,
    $I(Z_{1};C_{1})=I(Z_{2};C_{2})=0.$
@@ 3 Information-guided disentangled representation seeking (IndiSeek)
$\max_{Z_{i}}\;I(Z_{i},C_{i};X_{i})\qquad\text{subject to}\quad I(Z_{i};C_{i})\approx 0.$ (3)
This objective reflects the prevailing philosophy: modality-specific features should be maximally informative about their own modality, conditional on the shared component, while being independent of it.
Our position is not to redefine this conceptual objective, but to propose an effective way to implement it in practice. In the following subsections, we describe a two-stage strategy to this end: (1) learning shared features via CLIP, and (2) extracting modality-specific features with a reconstruction-guided objective conditional on learned shared features in step (1). See Figure 3 for an overview of IndiSeek.
@@ 3.3 Comparison with two SOTA Methods
Recent approaches to disentangled multi-modal representation learning differ mainly in how they approximate the two terms in Eq. (4): the disentanglement enforcing term $I(Z_{i},C_{i})$, and the complementary information capture term $I(Z_{i},C_{i};X_{i})$.
@@
liang2021multibench also proposes a self-supervised variant of Factorized CL by leveraging task-relevant augmentations, more discussions on which are deferred to Section 6 and Appendix A.1. In this paper, we adapt the unsupervised version of Factorized CL, i.e., we use CLUBInfoNCECritic between $Z_{1}$ and $C_{1}$ in liang2023factorizedgithub as the disentanglement objective, and adopt the implementation of CLUB loss in cheng2020club in both IndiSeek  and Factorized CL for fair comparison. (more details are presented in Appendix A.1).
@@
We note that, in the reconstruction term, wang2024information adopts the shared feature from the other modality, i.e., the actual reconstruction of interest becomes $I(Z_{i},C_{3-i};X_{i})$ instead of $I(Z_{i},C_{i};X_{i})$.
@@ 4 Experiments with a CITE-seq dataset
Figure 4: Performance of IndiSeek  in CITE-seq dataset ($\lambda=1$).
@@
In the left panel of Figure 4, $27$ level-2 cell types are ranked according to average $\theta(c)$ scores of cells in the test set with respective cell type annotations, in which a higher rank indicates a higher impact of the RNA-specific information in determining the cell type. Each bar is colored according to the coarser level-1 cell type that each level-2 cell type belongs to for easier visual inspection. Different cell types within the T cell family have negative scores, indicating RNA-specific information is less informative compared to ADT-specific counterpart for differentiating different T cell subpopulations, which is in line with the observation that clustering transcriptomics data typically does not delineate T cell subpopulations (szabo2019single; zheng2021pan) and gating based on surface protein markers is often needed (kotliar2025reproducible). In contrast, RNA-specific information plays a more important role in distinguishing different progenitor cells than ADT-specific information. To further demonstrate the quality of learned modality-specific representations, we plot the ranks of average $\theta(c)$ values against the benchmark ranks of average RNA weights in hao2021integrated of the 27 level-2 cell types in the right panel in Figure 4, which exhibit a strong correlation that is confirmed by Spearman’s rank correlation (spearman1961proof) of $0.910$. In Appendix A.3, we also present the UMAP of learned representations, which exhibits clear clusters with respect to two levels of cell types. As a comparison, we repeated the foregoing experiments across $10$ random seeds while replacing IndiSeek with InfoDisen or FactorizedCL when training the same neural net architecture and the resulting alignment metrics with respect to the results in hao2021integrated are presented in Figure 5, in which we also report the error bars for results over $10$ random seeds. Both alternative methods yielded inferior alignment with the relative importance of the RNA modality in hao2021integrated compared to IndiSeek. To ablate the choice of $\lambda$, in all experiments with real-world data, we report results with varying $\lambda$ on grid values $10^{j}$ for j=3,2,1,0,1,2,3j=-3,-2,-1,0,1,2,3.
@@ 5 Experiments with MultiBench datasets
Table 1: Comparison of accuracy on multimodal datasets (max over $\lambda$, averaged over 10 seeds). All values are percentages.
@@
Table 2: MOSI results with varying output dimensions (max over $\lambda$, averaged over 10 seeds). All values are percentages.
@@
Table 3: Multi-task performance on MIMIC dataset (ICD-9 groups 17–19). All values are percentages.
@@ 6 Extensions and discussion
The selection of the Lagrange multiplier $\lambda$ in (6) is important for the implementation of IndiSeek. With the rescaled reconstruction loss, the CLUB upper bound and the reconstruction term are now both scale-free, and our experimental results have suggested a candidate range of $\lambda$’s between $10^{-1}$ to $10^{1}$. Identifying the precise optimal choice of $\lambda$ remains an interesting future direction to enhance empirical performance and theoretical understanding.
@@ A.1 Implementation details and metrics
We follow the implementation of CLUB in cheng2020club. As CLUB loss is a variational upper bound for mutual information, we use an auxiliary conditional density $q_{\theta}(z|c)$ as an approximation for $p(c|z)$, where $q_{\theta}(z|c)$ lies in the Gaussian family, and we optimize the moments of the Gaussian distribution using 5-layer MLPs in each epoch. The inner optimization is run for 5 epochs at each epoch to optimize representations, and the implementation is the same for both IndiSeek (we use cross-modal shared features) and FactorCL for fair comparison.
For FactorizedCL, liang2021multibench also proposes a self-supervised variant of Factorized CL by leveraging task-relevant augmentations, more discussions on which are deferred to Section 6. In this paper, we adapt the unsupervised version of Factorized CL, i.e., we use CLUBInfoNCECritic between $Z_{1}$ and $Z_{2}$ in liang2023factorizedgithub as the disentanglement objective, and adopt the implementation of CLUB loss in cheng2020club in both IndiSeek  and Factorized CL for fair comparison.
@@ A.2 Numerical simulations
For Setting 1, given shared features are generated by sine and cubic functions, we can see from Figure 6 and 7 that both Factorized CL and IndiSeek are capable of extracting modality-specific features that are independent of shared ones when $\lambda=0.01$, while IndiSeek tend to be more robust to the choice of $\lambda$ and preserves the desired performance when $\lambda=0.1$. However, in contrast, InfoDisen, due to the limitation of orthogonal loss in handling nonlinear dependence, fails to disentangle modality-specific features from shared ones regardless of the value of $\lambda$.
@@
From Figures 9, 10, and 11, we can see that similar to the results presented in Section 1.1, with $\lambda$ in the range of $\{10^{j}:j=-2,-1,0,1\}$, Factorized CL fails to enforce independence between shared and modality-specific features; InfoDisen tends to miss modality-specific features with small $\lambda$ and obtain dependent features due to the limitations of orthogonal loss in handling nonlinear dependence. In comparison, with $\lambda=0.1$, IndiSeek can exactly capture the remaining two modality-specific features. Moreover, these figures also reveal the tradeoff between the disentanglement and sufficiency, determined by the value of $\lambda$. With $\lambda=0.01$, more weight is put on the disentanglement between shared and modality-specific features; thus, with a strong regularization on dependence, IndiSeek and InfoDisen tend to miss one modality-specific feature. On the other hand, with $\lambda=10$, the loss puts more weight on reconstruction/sufficiency, and as a result, disentangled methods may extract features that are dependent on the shared ones.
@@ A.3 CITE-seq dataset implementation details
We consider the bone marrow CITE-seq dataset from stuart2019comprehensive which consists of measurements in two modalities on $30672$ individual cells: transcriptome (RNA) and $25$ cell-surface proteins (ADT). We randomly split the dataset into a training set of size $15000$ and a test set of size $15672$. We follow the preprocessing in https://satijalab.org/seurat/articles/weighted_nearest_neighbor_analysis, where we normalize ADT data with centering to produce a 24-dimensional input based on moments estimated from the training set only, and normalize the extremely high-dimensional RNA data and extract the first 200 principal components as the inputs for CLIP.
In the experiments presented in Figure 4 and Table 5, architectures of representation maps $g$ and $h$ are all $5$-layer ReLU networks with width $50$ and output dimension $50$. Throughout experiments with real-world datasets, we adopt a batch size of $128$ and a learning rate of $10^{-4}$ for training, in which we set the number of epochs to $2000$.
In Figure 18, Figure 19, and Figure 20, we also visualize the learned representations $(C_{i},Z_{i})$ by all three methods with $\lambda=1.0$ for each modality. In the left panels, the representations are colored by level-1 cell types, and in the right panels, to ease visualization, we plot a subset of representations with level-2 cell types that are among the top-$40\%$ in the benchmark weights in the corresponding modality.
To quantify the separation of clusters, we adopt two metrics: (1) the Adjusted Rand Index (ARI) (hubert1985comparing), which measures agreement between $k$-means clustering on the learned representations and the ground-truth labels, with values ranging from 1-1 to $1$ where higher values indicate better cluster recovery; and (2) the Silhouette Score (rousseeuw1987silhouettes), which quantifies cluster compactness and separation using the ground-truth labels, with values in [1,1][-1,1] where higher scores indicate better-defined clusters.
@@
Figure 19: UMAP of FactorizedCL-learned representations $(C_{i},Z_{i})$ by IndiSeek (colored by cell types).
Figure 20: UMAP of InfoDisen-learned representations $(C_{i},Z_{i})$ by IndiSeek (colored by cell types).

### v4

@@ 1 Introduction
In contrast, identifying modality-specific features, also known as disentangled representation learning, has become an active research area (fischer2020conditional; liang2023factorized; liu2023focal; wang2024information; dufumier2024align; wang2024decoupling). While shared features should be minimally sufficient, modality-specific features must both remain independent of (i.e., disentangled from) shared features and capture all complementary information within each modality (liang2023factorized; wang2024information; wang2024decoupling). From an information-theoretic perspective, this requires (1) simultaneously minimizing the mutual information between shared and modality-specific features and (2) maximizing the mutual information between data and their union. Multiple state-of-the-art methods (liang2023factorized; wang2024information) tackle this problem from such a perspective, but differ in how they approximate the underlying mutual information. Since such quantities are difficult to estimate with finite samples, these methods often trade off independence against completeness. As we show in the motivating examples below, this tradeoff could interfere with their ability to achieve both goals simultaneously.
@@ 1.1 Motivating examples and limitations of SOTA
We present the average coordinate importance metrics in learned modality-specific representation maps over $50$ simulation runs in Figures 1 and 2 for three different learning algorithms in the two settings, respectively. In both figures, the left panels report results of the SOTA approach FactorizedCL in liang2023factorized (without self-supervision), the middle panels report results of the SOTA approach InfoDisen in wang2024information, and the right panels report results of the new approach IndiSeek we propose in this paper. Tuning parameters [^2] See Appendix A.2 for an ablation study on performances with different tuning parameter values. are set at $0.1$ for InfoDisen and IndiSeek and at $1.0$ by default for FactorizedCL.
As we have reasoned when introducing the settings, the ideal modality-specific representations in Setting 1 should depend on and only on the last four coordinates, while in Setting 2, on and only the last two coordinates. The left and the middle panels of both figures suggest that neither FactorizedCL nor InfoDisen learn an ideal representation in either setting, as irrelevant coordinates play non-trivial roles in determining the learned modality-specific representations. Therefore, both SOTA approaches have their limitations in achieving the desired goal of disentangled representation learning. In contrast, the right panel of Figure 1 shows that the new method IndiSeek we are to propose achieves the ideal property: the learned modality-specific features depend on and only on all coordinates that are independent of those involved in shared features. In Figure 2, compared with SOTA methods, modality-specific representations learned by IndiSeek exhibit much less dependence on both shared and redundant features, which demonstrates its effectiveness in this setting with highly nonlinear dependence and nontrivial redundancy.
@@ 2 Ideal decomposition of multi-modal data
    Independence: modality-specific and shared features should be disentangled [^4] As $C_{1}$ is restricted to be a function of $X_{1}$, it is possible that $C_{1}$ also contains modality-specific information. Meanwhile, all information in $C_{2}$ about $X_{1}$ is strictly shared. Therefore, we require $I(Z_{1};C_{2})=I(Z_{2};C_{1})=0$ instead of $I(Z_{1};C_{1})=I(Z_{2};C_{2})=0$.,
    $I(Z_{1};C_{2})=I(Z_{2};C_{1})=0.$
@@ 3 Information-guided disentangled representation seeking (IndiSeek)
$\max_{Z_{i}}\;I(Z_{i},C_{j};X_{i})\qquad\text{subject to}\quad I(Z_{i};C_{j^{\prime}})\approx 0,$ (3)
where $j,j^{\prime}\in\{1,2\}$ have respective specifications in each method, in which cross-modal switching of shared features is adopted when $j,j^{\prime}\neq i$. This objective reflects the prevailing philosophy: modality-specific features should be maximally informative about their own modality, conditional on the shared component, while being independent of it.
Our position is not only to refine this conceptual objective to relieve the issue with redundancy in shared features, but also to propose an effective and tractable way to implement it in practice. In the following subsections, we describe a two-stage strategy to this end: (1) learning shared features via CLIP, and (2) extracting modality-specific features with a reconstruction-guided objective conditional on learned shared features in step (1). See Figure 3 for an overview of IndiSeek.
@@ 3.3 Comparison with two SOTA Methods
Recent approaches to disentangled multi-modal representation learning differ mainly in how they approximate the two terms in Eq. (3): the disentanglement enforcing term $I(Z_{i},C_{j})$, and the complementary information capture term $I(Z_{i},C_{j^{\prime}};X_{i})$, where $j,j^{\prime}\in\{1,2\}$ will be specified in each method.
@@
We note that here $j,j^{\prime}$ in (3) both equal $i$, i.e., no cross-modal switching is implemented. liang2021multibench also proposes a self-supervised variant of Factorized CL by leveraging task-relevant augmentations, more discussions on which are deferred to Section 6 and Appendix A.1. In this paper, we adapt the unsupervised version of Factorized CL, i.e., we use CLUBInfoNCECritic between corresponding features in liang2023factorizedgithub as the disentanglement objective, and adopt the implementation of CLUB loss in cheng2020club in both IndiSeek  and Factorized CL for fair comparison. (more details are presented in Appendix A.1).
@@
We note that, in the reconstruction term, wang2024information adopts the shared feature from the other modality, i.e., the actual reconstruction of interest becomes $I(Z_{i},C_{3-i};X_{i})$ instead of $I(Z_{i},C_{i};X_{i})$ by replacing $j$ with 3i3-i in (3).
@@ 4 Experiments with a CITE-seq dataset
Figure 4: Performance of IndiSeek in CITE-seq dataset ($\lambda=10.0$).
@@
In the left panel of Figure 4, $27$ level-2 cell types are ranked according to average $\theta(c)$ scores of cells in the test set with respective cell type annotations, in which a higher rank indicates a higher impact of the RNA-specific information in determining the cell type. Each bar is colored according to the coarser level-1 cell type that each level-2 cell type belongs to for easier visual inspection. Different cell types within the T cell family have negative scores, indicating RNA-specific information is less informative compared to ADT-specific counterpart for differentiating different T cell subpopulations, which is in line with the observation that clustering transcriptomics data typically does not delineate T cell subpopulations (szabo2019single; zheng2021pan) and gating based on surface protein markers is often needed (kotliar2025reproducible). In contrast, RNA-specific information plays a more important role in distinguishing different progenitor cells than ADT-specific information. To further demonstrate the quality of learned modality-specific representations, we plot the ranks of average $\theta(c)$ values against the benchmark ranks of average RNA weights in hao2021integrated of the 27 level-2 cell types in the right panel in Figure 4, which exhibit a strong correlation that is confirmed by Spearman’s rank correlation (spearman1961proof) of $0.910$. In Appendix A.3, we present UMAPs of IndiSeek-learned representations in Figures 18 and 21, which exhibit clear clusters with respect to two levels of cell types that are also confirmed with metrics such as ARI and NMI (details are given in Appendix A.3). As a comparison, we repeated the foregoing experiments across $10$ random seeds while replacing IndiSeek with InfoDisen or FactorizedCL when training the same neural net architecture and the resulting alignment metrics with respect to the results in hao2021integrated are presented in Figure 5, in which we also report the error bars for results over $10$ random seeds. Both alternative methods yielded inferior alignment with the relative importance of the RNA modality in hao2021integrated compared to IndiSeek. To ablate the choice of $\lambda$, in all experiments with real-world data, we report results with varying $\lambda$ on grid values $10^{j}$ for j=3,2,1,0,1,2,3j=-3,-2,-1,0,1,2,3. Additionally, UMAPs of representations learned by InfoDisen and FactorizedCL are also presented in Appendix A.3 Figures 19, 20, 22, and 23.
@@ 5 Experiments with MultiBench datasets
Table 1: Comparison of accuracy on multimodal datasets (averaged over 10 seeds, standard errors in parentheses, max average over $\lambda$). All values are percentages.
@@
Table 2: MOSI results with varying output dimensions (averaged over 10 seeds, standard errors in parentheses, max average over $\lambda$). All values are percentages.
@@
Table 3: Multi-task performance (averaged over 10 seeds, standard errors in parentheses, max average over $\lambda$) on MIMIC dataset (ICD-9 groups 17–19). All values are percentages.
@@ 6 Extensions and discussion
The selection of the Lagrange multiplier $\lambda$ in (6) is important for the implementation of IndiSeek. With the rescaled reconstruction loss, both the CLUB upper bound and the reconstruction term are scale-free, and our experimental results have suggested a candidate range of $\lambda$’s between $10^{-1}$ to $10^{1}$. Identifying the precise optimal choice of $\lambda$ remains an interesting future direction to enhance empirical performance and theoretical understanding.
@@ A.1 Implementation details and metrics
We follow the implementation of CLUB in cheng2020club. As CLUB loss is a variational upper bound for mutual information, we use an auxiliary conditional density $q_{\theta}(z|c)$ as an approximation for $p(c|z)$, where $q_{\theta}(z|c)$ lies in the Gaussian family, and we optimize the moments of the Gaussian distribution using 5-layer MLPs in each epoch. The inner optimization is run for 5 epochs at each epoch to optimize representations, and the implementation the CLUB loss is the same for both IndiSeek and FactorizedCL for fair comparison.
liang2021multibench also proposes a self-supervised variant of Factorized CL by leveraging task-relevant augmentations, more discussions on which can be found in Section 6. In this paper, we adapt the unsupervised version of Factorized CL, i.e., we use CLUBInfoNCECritic between corresponding features in liang2023factorizedgithub as the disentanglement objective, and adopt the implementation of CLUB loss in cheng2020club in both IndiSeek  and Factorized CL for fair comparison.
@@ A.2 Numerical simulations
For Setting 1, given shared features are generated by sine and cubic functions, we can see from Figure 6 and 7 that both Factorized CL and IndiSeek are capable of extracting modality-specific features that are independent of shared ones when $\lambda=0.01$, while IndiSeek tend to be more robust to the choice of $\lambda$ and preserves the desired performance when $\lambda=0.1$. However, InfoDisen, due to the limitation of orthogonal loss in handling nonlinear dependence, fails to disentangle modality-specific features from shared ones regardless of the value of $\lambda$.
@@
From Figures 9, 10, and 11, we can see that similar to the results presented in Section 1.1, with $\lambda$ in the range of $\{10^{j}:j=-2,-1,0,1\}$, InfoDisen fails to enforce independence between shared and modality-specific features; FactorizedCL tends to exclude shared features with a small $\lambda$ but still captures at least one of redundant features that is correlated with shared ones. In comparison, with $\lambda=0.01$, IndiSeek can exactly capture the remaining two modality-specific features. Moreover, these figures also reveal the tradeoff between the disentanglement and sufficiency, determined by the value of $\lambda$. With $\lambda=0.01$, more weight is put on the disentanglement between shared and modality-specific features; thus, with a strong regularization on dependence, IndiSeek tends to exclude all shared and redundant features. On the other hand, with $\lambda=10$, the loss puts more weight on reconstruction/sufficiency, and as a result, disentangled methods may extract features that are dependent on the shared ones.
@@ A.3 CITE-seq dataset implementation details
We consider the bone marrow CITE-seq dataset from stuart2019comprehensive which consists of measurements in two modalities on $30672$ individual cells: transcriptome (RNA) and $25$ cell-surface proteins (ADT). We randomly split the dataset into a training set of size $15000$ and a test set of size $15672$. We follow the preprocessing in https://satijalab.org/seurat/articles/weighted_nearest_neighbor_analysis, where we normalize ADT data with centering to produce a 24-dimensional input based on moments estimated from the training set only, and normalize the high-dimensional RNA data and extract the first 200 principal components as the inputs for CLIP.
In the experiments presented in Figure 4 and Table 5, architectures of representation maps $g$ and $h$ for three methods are all $5$-layer ReLU networks with width $50$ and output dimension $50$. Throughout experiments with real-world datasets, we adopt a batch size of $128$ and a learning rate of $10^{-4}$ for training, in which we set the number of epochs to $2000$.
In Figure 18, Figure 19, and Figure 20, we visualize the learned representations $(C_{i},Z_{i})$ by all three methods for each modality. The top rows correspond to the RNA modality, and the bottom rows to the ADT modality. In the left panels, the representations are colored by level-1 cell types, and in the right panels, to ease visualization, we plot representations of a subset of cells whose level-2 cell types are among the top-$40\%$ in the average benchmark weights across all level-2 cell types in the corresponding modality. In other words, the chosen level-2 cell types for each modality are those for which the corresponding modality is more important for their distinction. If a disentangled learning approach works well, the resulting modality-specific representations of the chosen level-2 cell types are expected to be well separated in the respective UMAPs. Additionally, in Figure 21, Figure 22, and Figure 23, we visualize the concatenated representations $(C_{1},C_{2},Z_{1},Z_{2})$ learned by all three methods, in which representations are colored by level-1 and level-2 cell types in the left and the right panels, respectively. Here, we choose the optimal $\lambda$ implied by Figure 5 for each method, respectively: $0.1$ for FactorizedCL, $1.0$ for InfoDisen, and $10.0$ for IndiSeek.
To quantify the separation of cell types, we adopt two metrics: (1) the Adjusted Rand Index (ARI) (hubert1985comparing), which measures agreement between $k$-means clustering on the learned representations and the ground-truth labels, with values ranging from 1-1 to $1$ where higher values indicate better cluster recovery; and (2) NMI (normalized mutual information) following (luecken2022benchmarking; pedregosa2011scikit), which quantifies the mutual dependence between $k$-means clustering and the ground-truth labels, with values in $[0,1]$ where higher scores indicate better alignment between the discovered clusters and true labels. Figures 18-23 demonstrate that, when measured by these two metrics, IndiSeek outperforms the two SOTA methods in both learned modality-specific representations and the overall learned representations.
@@
Figure 19: UMAP of FactorizedCL-learned representations $(C_{i},Z_{i})$ (colored by cell types).
Figure 20: UMAP of InfoDisen-learned representations $(C_{i},Z_{i})$ (colored by cell types).
Figure 21: UMAP of concatenated IndiSeek-learned representations $(C_{1},C_{2},Z_{1},Z_{2})$ (colored by cell types).
Figure 22: UMAP of concatenated FactorizedCL-learned representations $(C_{1},C_{2},Z_{1},Z_{2})$ (colored by cell types).
Figure 23: UMAP of concatenated InfoDisen-learned representations $(C_{1},C_{2},Z_{1},Z_{2})$ (colored by cell types).
A.3.1 CITE-seq results without dimension reduction of RNA data
We also present the results with CITE-seq data following the preprocessing in https://satijalab.org/seurat/articles/weighted_nearest_neighbor_analysis, where we normalize ADT data with centering to produce a 24-dimensional input based on moments estimated from the training set only, and normalize the extremely high-dimensional RNA data. Here, instead of extracting the first 200 principal components as the inputs for CLIP, we adopt the raw normalized $2000$-dimensional RNA data for training. The results are presented in Figure 24 and Figure 25, respectively. Compared with Figures 4-5, results with and without the PCA dimension reduction for the RNA modality are comparable.
Figure 24: Performance of IndiSeek  in CITE-seq dataset ($\lambda=10.0$).